# Deep Emotion Recognition in Textual Conversations: A Survey

### v3

@@ Abstract
While Emotion Recognition in Conversations (ERC) has seen a tremendous advancement in the last few years, new applications and implementation scenarios present novel challenges and opportunities. These range from leveraging the conversational context, speaker and emotion dynamics modelling, to interpreting common sense expressions, informal language and sarcasm, addressing challenges of real time ERC, recognizing emotion causes, different taxonomies across datasets, multilingual ERC to interpretability. This survey starts by introducing ERC, elaborating on the challenges and opportunities pertaining to this task. It proceeds with a description of the emotion taxonomies and a variety of ERC benchmark datasets employing such taxonomies. This is followed by descriptions of the most prominent works in ERC with explanations of the Deep Learning architectures employed. Then, it provides advisable ERC practices towards better frameworks, elaborating on methods to deal with subjectivity in annotations and modelling and methods to deal with the typically unbalanced ERC datasets. Finally, it presents systematic review tables comparing several works regarding the methods used and their performance. The survey highlights the advantage of leveraging techniques to address unbalanced data, the exploration of mixed emotions and the benefits of incorporating annotation subjectivity in the learning phase.
@@ 1 Introduction
Emotion recognition capabilities are essential not only for successful interpersonal relationships, but also for human-machine interaction. Understanding and knowing how to react to emotions significantly improves the interaction and its outcome. It is therefore a crucial component in the development of empathetic machines, which substantially enriches the experiences these can provide.
Emotion Recognition in Conversation (ERC) modules are useful for a wide range of applications, from automatic opinion mining, to emotion-aware conversational agents and as assisting modules for therapeutic practices. ERC is therefore an actively growing research field, with high applicability and potential. Figure 1 depicts the evolution of the number of ERC publications across the years, using the most established benchmark datasets [8][97][139][70], supporting the observation that the field is indeed growing. Note that it does not account for datasets that contain emotion causes, a reseach direction that recently gained traction, neither datasets from shared tasks.
@@
It should be noted that emotions are only fully recognizable by the person having the emotion. Thus emotion recognition should be interpreted as "inferring an emotional state from observations of emotional expressions and behavior, and through reasoning about an emotion-generating situation" [93].
Progress in ERC benefits from the advancements in Sentiment Analysis, particularly in dialogues, whose tasks include determining the polarity and emotion of text. Research in Sentiment Analysis has come a long way and there is a perception that the field has reached its maturity, but there are still aspects and sub-tasks that need to be thoroughly addressed in order to attain true sentiment understanding [96]. Such statement applies to ERC as well. Emotion and sentiment are used interchangeably throughout this survey, but one should bear in mind they have different meanings: emotion is a psychological state while sentiment is the mental attitude created through the existence of the emotion [16].
The conversational setting presents additional challenges and opportunities, besides the ones pertaining to Emotion Recognition, such as the presence of intricate dependencies between the emotional states of the subjects who participate in the conversation, and the use of context and common sense to express emotions, amongst others described in this survey.
@@ 1.1 Survey Contributions
To the best of our knowledge, the existent surveys on Emotion Recognition in Conversations are by Poria et al [99], focusing on the challenges and advances in text ERC, pointing out conversational context modelling, speaker specific modelling, the presence of emotion shift, multiparty conversations and the presence of sarcasm, and a more recent survey by Fu et al [27], on multimodal ERC, that covers context modeling in conversations, speaker dependency, methods for fusing multimodal information and presents challenges such as real world ERC and classification on the typical ERC imbalanced datasets.
Our survey discusses additional challenges for this task, such as recognizing emotion causes, different taxonomies across datasets, multilingual ERC and interpretability. It also extensively covers solutions to dealing with the imbalanced ERC datasets and to learning with subjectivity from annotators.
This is also the first survey presenting an extensive list of Deep Learning works in ERC, stating their methods, modalities and performance across various datasets, while also describing key works on the topic.
Moreoever, our survey provides comprehensive descriptions of Deep Learning methods from the Multi Layer Perceptron, Recurrent Neural Networks, Long Short-Term Memory Networks, Gated Recurrent Units, Convolutional Neural Networks, Graph Neural Networks, Attention Mechanisms to the Transformer, in the light of Emotion Recognition in Conversations. While these methods are transversal to all emotion recognition tasks, and therefore all emotion recognition surveys, some are more appropriate for ERC. Examples are Transformer-based architectures that are better at preserving long-term dependencies between sequences of utterances than recurrent neural networks.
Related relevant surveys covered sentiment analysis research challenges and directions [96] and textual emotion recognition and its challenges [19], the latter including a section on textual emotion recognition in dialogue, elaborating on utterance context modelling and dynamic emotion modelling. However, these are not specific for conversations, not specifying how the models and practices employed in general textual emotion recognition relate to conversations, neither covering the several challenges we cover since such challenges pertain to ERC.
We also set up an online repository [^1] http://github.com/patricia-pereira/deep-emotion-recognition-in-textual-conversations with the sources we present, organized into categories such as overviews, datasets and works on ERC.
@@ 1.2 Survey Methodology
For the review of ERC works, we have used the keyword “emotion” for most of the research references in top NLP conferences (ACL, EMNLP, NAACL, COLING, Findings) and also other conferences such as AAAI, INTERSPEECH and ICASSP and journals such as IEEE Access and TAFFC. We also resorted to searching mechanisms that output similar papers when prompted with a paper such as Connected Papers. For the sections on specific topics, we searched for the respective topics’ keywords.
@@ 1.3 Outline
This survey is organised as follows: Section 1 introduces the topic of Emotion Recognition in Conversations and Section 2 presents its challenges and opportunities. Section 3 describes the emotion taxonomies and a variety of ERC benchmark datasets employing such taxonomies. Section 4 provides descriptions of the most prominent works in ERC and explains the Deep Learning architectures employed. Section 5 describes advisable ERC practices towards better frameworks, elaborating on methods to deal with subjectivity in annotations and methods to deal with the typically unbalanced ERC datasets. Section 6 consists in systematic review tables comparing several works regarding the methods used and their performance. Section 7 provides insights into how the aformentioned works attempted to solve the ERC challenges presented in section 2. Finally, Section 8 presents the conclusions and takeaways of the study.
@@ 2 ERC Challenges and Opportunities
Research in ERC has come a long way and several challenges have been addressed, as described in this section. From context, to speaker and emotion shift modelling to interpreting common sense, informal language and sarcasm, Deep Learning architectures have demonstrated its potential to address this aspects, although there is room for improvement. Application scenarios such as real time ERC pose additional challenges. In this section, we describe the challenges and point to the sections that are connected with each challenge and in Section 7 we wrap up the challenges, elaborating on how they are currently being addressed.
@@ 2.1 Context, Speaker and Emotion Dynamics Modelling
While early works [60][129] have tackled challenges from conversational context modelling, to speaker specific modelling and emotion dynamics, recent ERC approaches have leveraged the novel developments in Deep Learning to address this aspects.
Context modelling can be interpreted as leveraging information from previous or surrounding words or utterances. Speaker specific modelling leverages information about each speaker’s utterances, capitalizing on the fact that an utterance belongs to a given speaker, being extremely important in multiparty conversations. Emotion dynamics concerns the variations of the emotions of the participants in the conversation, being more informative but also harder to model than emotion shift. Emotions concern not only who expresses them but also the receiver. For example, more conscious and reflective people tend to be more sensitive regarding other’s emotions.
Concerning these challenges, gated and graph-based neural networks and transformer-based architectures are successful at modelling long range dependencies between words in utterances and relations between utterances. Key ERC works on Section 4 are described with regards to how these Deep Learning architectures were employed to address these challenges.
@@ 2.2 Common Sense, Informal Language and Sarcasm
Contextual embeddings from language models, such as BERT [20] and RoBERTA [73], are pre-trained with large amounts of data, which makes them suitable to deal with informal language, sarcasm and the use of common sense to express emotions, although there is room for improvement. Most of the ERC works featured in this survey make use of these contextual embeddings.
Regarding common sense, commonsense knowledge bases such as COMET [5] were developed to aid in leveraging the meaning behind phrases employing such knowledge. Three key ERC works described in Section 4 make use of knowledge bases. These knowledge bases, however, are not specific for emotions, which could be an extension to consider.
Concerning sarcasm, speaker specific modelling can aid in identifying the sarcastic participants in the conversation, to more accurately classify their utterances [99].
2.3 Real Time ERC
Systems to be deployed in a real life conversational scenarios may require real time emotion recognition [24], which is challenging not only for modelling and recognition but also in annotation because emotion expressions have to be simultaneously identified as its boundaries [93]. Datasets such as the ones described in subsubsection 3.2.7 feature annotations in continuous time, being more suitable for real time ERC.
@@ 2.4 Recognizing Emotion Causes
Recognizing emotion causes, or what triggered the emotions in the conversation, is also an only recently addressed challenge [98][121][50]. It involves resorting to datasets enriched with emotion causes, guiding classifiers to identify not only the emotion at stake but also what caused that emotion to arise. Addressing this challenge can greatly contribute towards more complete emotion understanding.
@@ 2.6 Multilingual ERC
Most of the benchmark datasets used to evaluate models and develop ERC applications are in English, as it can be observed in subsection 3.2. Translating the English conversations to the target application languages is not ideal, since some emotion can be lost in translation. The collection and annotation of multilingual datasets for ERC would benefit research in this direction. There are also dialogues that involve code-switching, i.e., featuring several languages in the same dialogue or even utterance, posing an additional challenge [56].
@@ 2.7 Interpretability
A property lacking in most of the works resorting to deep learning architectures is interpretability. It is difficult to interpret the reasoning behind these models’ decisions and identify what are the characteristic features of each class. Examining attention weights are a way of determining to which features the models were attending to while computing classifications which can aid interpretability although there are some issues raised in the community [49, 126].
@@ 3 ERC Taxonomies and Benchmark Datasets
In this section, a description of different taxonomies, i.e., ways of representing emotions, is provided, followed by the presentation of various datasets that resort to such taxonomies. The "dialogues in the datasets reflect our daily communication way" [70], therefore encompassing several emotions across different taxonomies.
@@ 3.1 Emotion Taxonomies
One of the most important issues in the study of emotion is whether there are basic types of affect or whether affective states should be modelled as combinations of locations on shared dimensions [138]. These two dominant views corresponded to the categorical approach, defended by Panksepp [88], and to the dimensional approach, proposed by Russell [106], respectively. This is reflected in two main taxonomy approaches, that are described and compared in this section. The Plutchik [94] wheel of emotions is also described.
@@ 3.1.1 Dimensional vs Categorical Approach
The description of emotions can follow a dimensional approach, a continuous model that allows for an unlimited states’ description. Following this approach, an emotion is described by a range of values for valence, arousal and dominance, the so called VAD model [82]. Valence is a measure of pleasure or displeasure of an emotion: happiness has a positive valence, fear has a negative valence; arousal ranges from excitement to calmness: anger is high in arousal, sadness is low in arousal; dominance measures how much choice one has over an emotion: fear is low dominance, admiration is high dominance. The distribution of common emotions over the VAD 3-dimensional space can be observed in Figure 2.
@@
Although there are only three dimensions, the unlimited states’ possibility makes annotations hard.
A motivation for the use of the dimensional description could be the existence of mixed emotions, elaborated in subsection 3.1.3.
The description can also follow a categorical approach, in which an emotion is assigned to one or more labels from a set of classes. Comparing to the dimensional approach, this categorical discrete framework makes it easier to annotate emotions but is more limited. Several scholars proposed to identify a set of primary emotions. Ekman’s basic emotions constitute a standard for this emotion description, categorized into joy, anger, sadness, surprise, fear, and disgust [22].
@@ 3.1.2 Plutchik Wheel of Emotions
The Plutchik wheel of emotions, depicted in Figure 3, identifies eight core emotions, organized in opposite pairs of a wheel: sadness and joy, anger and fear, expectation and surprise, and trust and disgust. These have associated emotions according to its intensities and also combine to form new emotional states [94].

### v4

@@ Abstract
Emotion Recognition in Conversations (ERC) is a key step towards successful human-machine interaction. While the field has seen tremendous advancement in the last few years, new applications and implementation scenarios present novel challenges and opportunities. These range from leveraging the conversational context, speaker, and emotion dynamics modelling, to interpreting common sense expressions, informal language, and sarcasm, addressing challenges of real-time ERC, recognizing emotion causes, different taxonomies across datasets, multilingual ERC, and interpretability. This survey starts by introducing ERC, elaborating on the challenges and opportunities of this task. It proceeds with a description of the emotion taxonomies and a variety of ERC benchmark datasets employing such taxonomies. This is followed by descriptions comparing the most prominent works in ERC with explanations of the neural architectures employed. Then, it provides advisable ERC practices towards better frameworks, elaborating on methods to deal with subjectivity in annotations and modelling and methods to deal with the typically unbalanced ERC datasets. Finally, it presents systematic review tables comparing several works regarding the methods used and their performance. Benchmarking these works highlights resorting to pre-trained Transformer Language Models to extract utterance representations, using Gated and Graph Neural Networks to model the interactions between these utterances, and leveraging Generative Large Language Models to tackle ERC within a generative framework. This survey emphasizes the advantage of leveraging techniques to address unbalanced data, the exploration of mixed emotions, and the benefits of incorporating annotation subjectivity in the learning phase.
@@ 1 Introduction
This survey systematically reviews Deep Emotion Recognition in Conversations (ERC) works. It elaborates on several ERC challenges and opportunities, explaining how these are currently tackled by the reviewed works and discussing how these can be further addressed.
Emotion recognition capabilities are essential not only for successful interpersonal relationships but also for human-machine interaction. Understanding and knowing how to react to emotions significantly improves the interaction and its outcome. It is therefore a crucial component in the development of empathetic machines, which substantially enriches the experiences these can provide.
ERC modules are useful for a wide range of applications, from automatic opinion mining, to emotion-aware conversational agents and as assisting modules for therapeutic practices. ERC is therefore an actively growing research field, with high applicability and potential. Figure 1 depicts the evolution of the number of ERC publications across the years, using the most established benchmark datasets [6, 88, 135, 62], supporting the previous observation. Note that it does not account for datasets addressing emotion causes, a recent research direction.
@@
It should be noted that emotions are only fully recognizable by the person having the emotion. Thus emotion recognition should be interpreted as “inferring an emotional state from observations of emotional expressions and behavior, and through reasoning about an emotion-generating situation” [84].
Progress in ERC benefits from the advancements in Sentiment Analysis, particularly in dialogues, whose tasks include determining the polarity and emotion of text. Research in Sentiment Analysis has come a long way and there is a perception that the field has reached its maturity, but there are still aspects and sub-tasks that need to be thoroughly addressed to attain true sentiment understanding [87]. Such a statement applies to ERC as well. Emotion and sentiment are used interchangeably throughout this survey, but one should bear in mind they have different meanings: emotion is a psychological state while sentiment is the mental attitude created through the existence of the emotion [12].
The conversational setting presents additional challenges and opportunities, besides the ones of Emotion Recognition, such as the presence of intricate dependencies between the emotional states of the subjects who participate in the conversation, and the use of context and common sense to express emotions, amongst others described in this survey.
1.1 Task Definition
The task of ERC consists of determining the emotion of each utterance in a textual conversation, for which speaker information is provided for each utterance. Formally, given a sequence of $N$ utterances $[(u_{1},p_{1}),(u_{2},p_{2}),\dots,(u_{N},p_{N})]$, where each utterance $u_{i}$ is spoken by participant $p_{i}$, the goal is to predict the emotion label $e_{i}$ for each utterance $u_{i}$.
@@ 1.2 Survey Contributions
To the best of our knowledge, the existent surveys on Emotion Recognition in Conversations are by Poria et al [90], focusing on the challenges and advances in text ERC, pointing out conversational context modelling, speaker-specific modelling, the presence of emotion shift, multiparty conversations and the presence of sarcasm, and a more recent survey by Fu et al [23], on multimodal ERC, that covers context modeling in conversations, speaker dependency, methods for fusing multimodal information and presents challenges such as real-world ERC and classification on the typical ERC imbalanced datasets.
Related relevant surveys covered sentiment analysis research challenges and directions [87] and textual emotion recognition and its challenges [14], the latter including a section on textual emotion recognition in dialogue, elaborating on utterance context modelling and dynamic emotion modelling. However, these are not specific to conversations, not specifying how the models and practices employed in general textual emotion recognition relate to conversations, nor covering several ERC pertaining challenges.
Our survey discusses additional challenges for this task, such as recognizing emotion causes, different taxonomies across datasets, multilingual ERC, and interpretability. It also extensively covers solutions to dealing with imbalanced ERC datasets and learning with subjectivity from annotators.
This is also the first survey presenting an extensive list of Deep Learning works in ERC, stating their methods, modalities, and performance across various datasets, while also describing key works on the topic. We provide comprehensive descriptions of Deep Learning methods from the Multi-Layer Perceptron, Recurrent Neural Networks, Long Short-Term Memory Networks, Gated Recurrent Units, Convolutional Neural Networks, Graph Neural Networks, and Attention Mechanisms to the Transformer, in the light of Emotion Recognition in Conversations. While these methods are transversal to all emotion recognition tasks, and therefore all emotion recognition surveys, some are more appropriate for ERC. Examples are Transformer-based architectures that are better at preserving long-term dependencies between sequences of utterances than recurrent neural networks.
We also set up an online repository [^1] http://github.com/patricia-pereira/deep-emotion-recognition-in-textual-conversations with the sources we present, organized into categories such as overviews, datasets, and works on ERC.
@@ 1.3 Survey Methodology
We report the steps of our systematic review: identification, screening, eligibility, and inclusion.
In the identification step, to collect ERC works for Sections 4 and 6, we search the accepted paper’s sections of leading NLP conferences (ACL, EMNLP, NAACL, EACL, COLING, and Findings) and additional conferences (AAAI, INTERSPEECH, and ICASSP), using the title keywords “emotion” in combination with “conversational” or “conversations”. Journal papers are also collected from IEEE Access and IEEE Transactions on Affective Computing.
We limit our search to papers published after 2017, given the significant advances in Deep Learning models since that time, particularly the introduction of the Transformer architecture [112] and pre-trained Transformer models such as BERT [15] and RoBERTa [65].
In the screening step, the inclusion criteria are papers that address emotion recognition within conversations, report performance on one or more of the four main benchmark datasets listed in subsection 3.2 (IEMOCAP, MELD, EmoryNLP, and DailyDialog), and report F1-scores, the adequate evaluation metric for emotion recognition in conversations, as elaborated in subsection 5.2. The exclusion criteria are papers not reporting results on the aforementioned benchmark datasets or not reporting F1-scores.
For Sections 2, 3, and 5, we conduct topic-specific searches using Google Scholar, employing keywords such as “challenges in ERC” or “data taxonomies in emotion recognition”. Search results are filtered by relevance and publication date to prioritize recent and pertinent research.
To complement our searches, we use tools such as Connected Papers to discover papers related to the ones already retrieved.
To assess paper eligibility we evaluate each paper’s full text, focusing on the reported results and alignment with the survey scope, more specifically, the use of Deep Learning methods.
@@ 1.4 Outline
This survey is organised as follows: Section 1 introduces the topic of Emotion Recognition in Conversations and Section 2 presents its challenges and opportunities. Section 3 describes the emotion taxonomies and a variety of ERC benchmark datasets employing such taxonomies. Section 4 provides descriptions of the most prominent works in ERC and explains the Deep Learning architectures employed. Section 5 describes advisable ERC practices towards better frameworks, elaborating on methods to deal with subjectivity in annotations and methods to deal with the typically unbalanced ERC datasets. Section 6 consists of systematic review tables comparing several works regarding the methods used and their performance and provides insights into how the aforementioned works attempted to solve the ERC challenges presented in section 2. Section 7 elaborates on future work directions. Finally, Section 8 presents the conclusions, limitations and key findings of the survey.
@@ 2 ERC Challenges and Opportunities
Research in ERC has come a long way and several challenges have been addressed, as described in this section. From context, to speaker and emotion shift modelling to interpreting common sense, informal language and sarcasm, Deep Learning architectures have demonstrated their potential to address these aspects, although there is room for improvement. Application scenarios such as real-time ERC pose additional challenges. In this section, we describe such challenges and point to the sections that are connected with each challenge, and in subsection 6.1 we wrap up the challenges, elaborating on how they are currently being addressed. In Section 7, we suggest future work directions concerning partly unaddressed challenges.
@@ 2.1 Context, Speaker and Emotion Dynamics Modelling
While early works [53, 120] have tackled challenges from conversational context modelling, to speaker-specific modelling and emotion dynamics, resorting to, for example, gated networks, recent ERC approaches have leveraged the novel developments in Deep Learning, such as Transformers and Transformer-based Pre-trained Models to address these aspects.
Context modelling can be interpreted as leveraging information from previous or surrounding words or utterances. Speaker-specific modelling leverages information about each speaker’s utterances, capitalizing on the fact that an utterance belongs to a given speaker, being extremely important in multiparty conversations. Emotion dynamics concerns the variations of the emotions of the participants in the conversation, being more informative but also harder to model than emotion shift. Emotions concern not only who express them but also the receiver. For example, more conscious and reflective people tend to be more sensitive regarding other’s emotions.
Concerning these challenges, gated and graph-based neural networks and transformer-based architectures are successful at modelling long range dependencies between words in utterances and relations between utterances. Key ERC works in Section 4 are described with regards to how these Deep Learning architectures were employed to address these challenges.
@@ 2.2 Common Sense, Informal Language and Sarcasm
Contextual embeddings from language models, such as BERT [15] and RoBERTA [65], are pre-trained with large amounts of data, which makes them suitable to deal with informal language, sarcasm, and use of common sense to express emotions, although there is room for improvement. Most of the ERC works featured in this survey make use of these contextual embeddings.
Regarding common sense, commonsense knowledge bases such as COMET [4] were developed to aid in leveraging the meaning behind phrases employing such knowledge. Three key ERC works described in Section 4 make use of knowledge bases. These knowledge bases, however, are not specific to emotions, which could be an extension to consider.
Concerning sarcasm, speaker-specific modelling can aid in identifying the sarcastic participants in the conversation, to more accurately classify their utterances [90].
2.3 Real-Time ERC
Systems to be deployed in real-life conversational scenarios may require real-time emotion recognition [20], which is challenging not only for modelling and recognition but also in annotation because the boundaries of each expression of emotion have to be simultaneously recognized as the emotion category [84]. Datasets such as the ones described in subsubsection 3.2.5 feature annotations in continuous time, being more suitable for real-time ERC.
@@ 2.4 Recognizing Emotion Causes
Recognizing emotion causes, or what triggered the emotions in the conversation, is also an only recently addressed challenge [89, 113, 43, 38, 40, 68]. It involves resorting to datasets enriched with emotion causes, guiding classifiers to identify not only the emotion at stake but also what caused that emotion to arise. Addressing this challenge can greatly contribute towards a more complete emotional understanding.
@@ 2.6 Multilingual ERC
Most of the benchmark datasets used to evaluate models and develop ERC applications are in English, as it can be observed in subsection 3.2. Translating the English conversations to the target application languages is not ideal since some emotion can be lost in translation. The collection and annotation of multilingual datasets for ERC would benefit research in this direction. Some dialogues involve code-switching, i.e., featuring several languages in the same dialogue or even utterance [49]. The cultural aspect of emotion expression/interpretation also poses a challenge, since emotion is not universally expressed or understood in the same way across different cultures.
@@ 2.7 Interpretability
A property lacking in most of the works resorting to deep learning architectures is interpretability. It is difficult to interpret the reasoning behind these models’ decisions and identify what are the characteristic features of each class. Examining attention weights is a way of determining to which features the models were attending to while computing classifications, which can aid interpretability, although there are some issues raised in the community [42, 118]. Other methods can be proposed to provide more interpretability to the ERC pipeline, such as Fuzzy Fingerprinting Transformer Embedding Representations [82].
@@ 3 ERC Taxonomies and Benchmark Datasets
In this section, a description of different taxonomies, i.e., ways of representing emotions, is provided, followed by the presentation of various datasets that resort to such taxonomies. The “dialogues in the datasets reflect our daily communication way” [62], therefore encompassing several emotions across different taxonomies.
@@ 3.1 Emotion Taxonomies
One of the most important issues in the study of emotion is whether there are basic types of affect or whether affective states should be modelled as combinations of locations on shared dimensions [134]. These two dominant views corresponded to the categorical approach, defended by Panksepp [78], and to the dimensional approach, proposed by Russell [97], respectively. This is reflected in two main taxonomy approaches, that are described and compared in this section. The Plutchik wheel of emotions [85] is also described.
@@ 3.1.1 Dimensional vs Categorical Approach
The description of emotions can follow a dimensional approach, a continuous model that allows for an unlimited states’ description. Following this approach, emotion is described by a range of values for valence, arousal, and dominance, the so-called VAD model [73]. Valence is a measure of pleasure or displeasure of emotion: happiness has a positive valence, fear has a negative valence; arousal ranges from excitement to calmness: anger is high in arousal, sadness is low in arousal; dominance measures how much choice one has over an emotion: fear is low dominance, admiration is high dominance. The distribution of common emotions over the VAD 3-dimensional space can be observed in Figure 2.
@@
A motivation for the use of the dimensional description could be the existence of mixed emotions, elaborated in subsubsection 3.1.3.
The description can also follow a categorical approach, in which an emotion is assigned to one or more labels from a set of classes. Compared to the dimensional approach, this categorical discrete framework makes it easier to annotate emotions but is more limited. Several scholars proposed to identify a set of primary emotions. The basic emotions proposed by Ekman [18] constitute a standard for this emotion description, categorised into joy, anger, sadness, surprise, fear, and disgust.
@@ 3.1.2 Plutchik Wheel of Emotions
The Plutchik wheel of emotions, depicted in Figure 3, identifies eight core emotions, organised in opposite pairs of a wheel: sadness and joy, anger and fear, expectation and surprise, and trust and disgust. These have associated emotions according to their intensities and also combine to form new emotional states [85].